OpenStudy (anonymous):

What are the solutions of x2 + 5x − 24 = 0?

OpenStudy (zehanz):

Can you factor it first?

OpenStudy (anonymous):

idk

OpenStudy (zehanz):

It will be like this: (x+a)(x+b). To find a and b, you have to know that: a+b = 5 (the number of x terms) a*b = -24 (the constant at the end) E.g. a=3, b=-8. Then 3*(-8)=-24 YES!! But: 3 + -8 = -5 NO!! So 3 and -8 are not what we are looking for, but: we're not far off. Just search a little more!

OpenStudy (anonymous):

ightt thanks you

OpenStudy (zehanz):

YW!